=== 2019/1/30(Wed.) ===

=== Weekend ===

=== Schedule ===
  * Current status as of 1/22: IR and Green reached X end PD.
  * --(1/22 Surface inspection for PR mirrors.)-- Cleaning will be done after 1/31.
  * --(1/22 or 23 Surface inspection for SR mirrors.)--
  * --(1/24-25 Cleaning HR of SR3, SRM.)--

  * --(1/23 BS is aligned to Y-end PD by IR.)--
  * --(1/24-25 ITMY is aligned to REFL by IR.)--
  * --(1/24-25 ITMX is aligned to REFL by IR.)--
  * --(1/24-25 Beam position of IR at SR3 is confirmed how far from the center.)--

  * --(1/28-29 SR3 is aligned to SR2 center by IR.)--
  * --(1/30 Green beam is injected from SR2 AR surface to Y end PD.)--

  * 1/31-2/1 Removing bellows at PR and SR area.
  * 1/31-2/1 Cleaning HR of PR3 and HR, AR of SR2

  * 2/4-8 --(Alignment of SR2 -> SRM by IR)--, SRM -> OMMT by IR -> OMMT -> OSTM by IR, OSTM -> OMC by IR
  * 2/4-8 Alignment of ETMY -> POS by Green

  * 2/12-15 Closing door for central area, pumping down
   * needs drift monitor for VIS

==== Report of work (*:responsible person) ====
 * COM: Green laser reached to ETMY.
 * Pedestal was replaced.
 * ISS: PD was installed on second loop.
 * CAL: Green beam monitor at ETMY will work, and needs fine alignment for finalization.
 * ETMY lock was released, but need little more time to make it work.

==== Work plan (*:responsible person) ====
 * SRM: needs some fine alignment.
 * DGS: weekly maintenance from 9am to 1pm.
 * PEM: Channel names will be replaced on Monday.
 * ETMY should be ready by the next week for Pcal finalization and ETM alignment to BS.